# DESI constraints on $\alpha$-attractor inflationary models

### v2

@@ V.1 Inflation and post-inflation evolution
In order to explain the existence of DE today, the potential needs to account for the large mismatch between the inflationary plateau and the tiny energy scale of the DE. As a consequence, a crucial requirement for the scalar potential is to have a rapid fall at the end of inflation where the Universe undergoes a period of kination-dominated phase and the Hubble rate evolves as $H\propto a^{-3}$. Reheating may undergo during this phase and the Universe eventually enters to radiation-dominated phase Ford (1987); Chun et al. (2009), followed by the matter-dominated phase as in the standard big-bang model. Then, at later times the energy density of scalar field can dominate the Universe again, which accounts for the present DE.
In summary, in this unified scenario of inflation and DE, the Universe undergoes dominations of inflaton, kination, radiation, matter, and DE. The duration of inflation, which is parametrized by the e-folding number $N_{\ast}$, can be related to the post-inflationary evolution using the following equation
@@
-- $\frac{1}{12}\ln g_{*,{\rm reh}}+\frac{1-3w_{\rm reh}}{12(1+3w_{\rm reh})}\ln\left(\frac{\rho_{\rm reh}}{\rho_{\rm end}}\right)\,,$
@@
Note that the lower and upper bounds are far beyond the other theoretical limits. The lower bound on $T_{\rm reh}$ comes from BBN, which necessitates the completion of reheating before BBN, providing a much tighter lower bound of $T_{\rm reh}\gtrsim 10^{-3}$ GeV. The upper bound arises from the requirement that the reheating temperature must not exceed the inflation scale. Using the $1\sigma$ upper bound on $V_{\ast}^{1/4}/M_{\rm Pl}<0.00406$, we find $T_{\rm reh}<4.1\times 10^{15}$eV. Therefore, the current bounds from CMB and LSS data are too weak to constrain the reheating temperature effectively. Conversely, these theoretical limits can be used to tighten the bounds on model parameters by combining them with the CMB and LSS data. In the next subsection, we discuss such additional constraints from the theoretical bounds $T_{\rm reh}/{\rm GeV}=[10^{-3},4.1\times 10^{15}]$, as well as constraints obtained from the inflationary GW background.
@@ V.3 Combined constraints
Figure 4: The spectrum of inflationary GWs predicted in the $\alpha$-attractor model with an early kination phase is compared with the sensitivities of current (solid line with gray shading) and future GW experiments (dashed line). The red shaded region indicates the possible range of the GW spectrum, with the upper and lower curves corresponding to the range of reheating temperatures, $T_{\rm reh}/{\rm GeV}=[10^{-3},4.1\times 10^{15}]$, which mark the onset of BBN and the end of inflation (instant reheating).
@@
Figure 5: The constraint in the $\alpha$$n_{s}$ plane, obtained by CMB+Lensing, DESI and Pantheon+ likelihoods (blue contour, same as in Fig. 2), is plotted alongside the bounds from the GW amplitude at high frequencies and the theoretical limits on reheating temperature. The orange shaded area indicates the parameter region constrained by the BBN bound, $\Omega_{\rm GW}h^{2}<7.6\times 10^{-7}$. The red band indicates the region predicting a reasonable value of the reheating temperature, $T_{\rm reh}/{\rm GeV}=[10^{-3},4.1\times 10^{15}]$.
@@
By using the indicated range obtained from the MCMC analysis, $V_{\ast}^{1/4}/M_{\rm Pl}=[0.00318,0.00406]$ and $T_{\rm reh}/{\rm GeV}=[10^{-3},4.1\times 10^{15}]$, discussed in Sec. V.1, in Fig. 4, we show the GW spectrum using the formulas described in Sec. V.2. Notably, when the reheating temperature is low $T_{\rm reh}\lesssim 10^{5}$eV (corresponding to large e-folding number, $N_{\ast}\gtrsim 71$), the model predicts too large GW amplitude at high frequencies, which violates constraints from the BBN and LIGO-Virgo-KAGRA (LVK) O3 upper bound Abbott et al. (2021), shown by shaded gray region. Here, the BBN bound is based on the requirement that the energy density of GWs should not be too large, as it would alter the Hubble expansion rate and disrupt the successful BBN. The recent joint CMB+BBN analysis implies $\Omega_{\rm GW}h^{2}<1.3\times 10^{-6}$ at $2\sigma$ for $f>2\times 10^{-11}$z Yeh et al. (2022).
@@
This highlights the interesting fact that certain parameter regions can be further constrained by GW constraints. In Fig. 5, the parameter space allowed by the BBN bound $\Omega_{\rm GW}h^{2}<1.3\times 10^{-6}$ is shaded with orange. As indicated in Fig 4, the parameter space predicting too low reheating temperature (larger e-folding number, larger $n_{s}$) is excluded as it predicts too large GW spectrum at high frequencies. Furthermore, the red shaded region indicates where the parameters predict a reasonable reheating temperature. A too low reheating temperature ($T_{\rm reh}\lesssim 10^{-3}$eV) disrupts the successful BBN, while an excessively high reheating temperature ($T_{\rm reh}\sim 4.1\times 10^{15}$eV) implies instant reheating [^3] Note that in the context of supergravity, considerations on the overproduction of gravitinos can further constrain the reheating temperature to $T_{\rm reh}\lesssim 10^{8}-10^{9}$eV Kawasaki et al. (2006); Dimopoulos and Owen (2017) or even down to $10^{6}$eV Kohri et al. (2006).. This temperature corresponds to the e-folding number of $N_{\ast}\sim 67$ and when the e-folding is lower than it we need a phase with $w<1/3$ such as a early matter domination, which requires extra physics to achieve it in the picture of $\alpha$-attractor model. Therefore, the overlap of the red and orange shaded areas represents the allowed region, enabling us to reject more parameter space of the model.
@@ VI Conclusions
In this work, we explored the compatibility of a specific class of $\alpha$-attractor models when using the latest observational data, motivated by recent implications of evolving DE equation of state parameter from DESI. The $\alpha$-attractor models present a compelling and unified framework, motivated by high-energy physics and supergravity, as an alternative to traditional single-field slow-roll inflation. These models connect early- and late-time physics, offering a viable and physically motivated alternative to ad-hoc phenomenological models.
@@
We also considered a connection of the $\alpha$-attractor models with GWs by comparing the spectrum of inflationary GWs predicted in the $\alpha$-attractor model with early kination phase; see Fig. 4, compared with sensitivities of current/future GW experiments. We find the interesting fact that certain parameter regions can be further constrained by the BBN constraint on the GW amplitude, see Fig. 5. These constraints could be strengthened by future CMB B-mode measurements as well as constraints by interferometer experiments. Additionally, by considering an appropriate range of reheating temperatures, we demonstrated that the parameter space can be further tightened. However, the analysis we performed here is only indicative, thus we leave a more complete study for future work.
@@ Acknowledgements
The authors thank Nils Schöneberg for useful discussions and they acknowledge the use of the Finis Terrae III supercomputer, which is part of the Centro de Supercomputacion de Galicia (CESGA) and is funded by the Ministry of Science and Innovation, Xunta de Galicia and ERDF (European Regional Development Fund). The authors also acknowledge support from the research project PID2021-123012NB-C43 and the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033. GA’s and SK’s research is supported by the Spanish Attraccion de Talento contract no. 2019-T1/TIC-13177 granted by the Comunidad de Madrid. SK is partly supported by the I+D grant PID2020-118159GA-C42 funded by MCIN/AEI/10.13039/501100011033, the Consolidación Investigadora 2022 grant CNS2022-135211, and Japan Society for the Promotion of Science (JSPS) KAKENHI Grant no. 20H01899, 20H05853, and 23H00110. MC acknowledges support from the “Ramón Areces” Foundation through the “Programa de Ayudas Fundación Ramón Areces para la realización de Tesis Doctorales en Ciencias de la Vida y de la Materia 2023”.
@@ Appendix A Quantitative discussion of the $\alpha$-attractor models
The main effect of $\alpha$-attractor models at late times, compared to $\Lambda$CDM, is a change at the background expansion of the Universe causing deviations at a) the CMB spectra at low multipoles via the Integrated Sachs-Wolfe (ISW) effect, and b) at the angular diameter distance, affecting dircectly the BAO measurements.

### v3

@@ V.1 Inflation and post-inflation evolution
In order to explain the existence of DE today, the potential needs to account for the large mismatch between the inflationary plateau and the tiny energy scale of the DE. As a consequence, a crucial requirement for the scalar potential is to have a rapid fall at the end of inflation where the Universe undergoes a period of kination-dominated phase and the Hubble rate evolves as $H\propto a^{-3}$. Reheating may happen during this phase and the Universe eventually enters to radiation-dominated phase Ford (1987); Chun et al. (2009), followed by the matter-dominated phase as in the standard big-bang model. Then, at later times the energy density of scalar field can dominate the Universe again, which accounts for the present DE.
In summary, in this unified scenario of inflation and DE, the Universe undergoes periods dominated by inflaton, kination, radiation, matter, and DE. The duration of inflation, which is parametrized by the e-folding number $N_{\ast}$, can be related to the post-inflationary evolution using the following equation
@@
-- $\frac{1}{12}\ln g_{*,{\rm reh}}+\frac{1-3w_{\rm reh}}{12(1+w_{\rm reh})}\ln\left(\frac{\rho_{\rm reh}}{\rho_{\rm end}}\right)\,,$
@@
Note that the lower bound is far beyond the other theoretical limit. While the value of $T_{\rm reh}$ depends heavily on the underlying particle physics model, and no robust prediction exists, there are two key theoretical constraints to consider. The lower bound on $T_{\rm reh}$ comes from BBN, which necessitates the completion of reheating before BBN, providing a much tighter lower bound of $T_{\rm reh}\gtrsim 10^{-3}$ GeV. The upper bound arises from the requirement that the reheating temperature must not exceed the inflation scale. Using the $1\sigma$ upper bound on $V_{\ast}^{1/4}/M_{\rm Pl}<0.00403$, we find $T_{\rm reh}<4.0\times 10^{15}$eV. Therefore, the current bounds from CMB and LSS data do not establish a competitive lower bound; however, they do provide a relatively tight upper bound when compared to the theoretical limits. In the following subsections, we will consider the range of reheating temperature $T_{\rm reh}/{\rm GeV}=[10^{-3},1.7\times 10^{9}]$, where lower bound is determined by the BBN, while upper bound is derived from our CMB and LSS analysis. Then we discuss the implications of these observational bounds on the inflationary gravitational wave background. We will also demonstrate how the upper limit on high-frequency GWs, derived from BBN, can further constrain the parameter space of this model.
@@ V.3 Combined constraints
[…]
@@
Figure 5: The constraint in the $\alpha$$n_{s}$ plane, obtained by CMB+Lensing, DESI and Pantheon+ likelihoods (blue contour, same as in Fig. 2), is plotted alongside the bounds from the GW amplitude at high frequencies and the theoretical limits on reheating temperature. The orange shaded area indicates the parameter region constrained by the BBN bound, $\Omega_{\rm GW}h^{2}<1.3\times 10^{-6}$. The red band indicates the region predicting a reasonable value of the reheating temperature, $T_{\rm reh}/{\rm GeV}=[10^{-3},4.0\times 10^{15}]$.
@@
By using the indicated range obtained from the MCMC analysis, $V_{\ast}^{1/4}/M_{\rm Pl}=[0.00318,0.00406]$ and $T_{\rm reh}/{\rm GeV}=[10^{-3},1.7\times 10^{9}]$, discussed in Sec. V.1, in Fig. 4, we show the GW spectrum using the formulas described in Sec. V.2.  [^3] It is worth noting that there is also flexibility to adjust the parameters $w_{\rm reh}$ and $g_{*,{\rm reh}}$, both of which influences the GW spectral amplitude. The values adopted here $w_{\rm reh}=1$ (the value of kination domination) and $g_{*,{\rm reh}}=106.75$ (the value from the standard model) are well-motivated, but in principle, they could be set as free parameters. Notably, when the reheating temperature is low $T_{\rm reh}\lesssim 10^{5}$eV (corresponding to large e-folding number, $N_{\ast}\gtrsim 64$), the model predicts too large GW amplitude at high frequencies, which violates constraints from the BBN and LIGO-Virgo-KAGRA (LVK) O3 upper bound Abbott et al. (2021), shown by shaded gray region. Here, the BBN bound is based on the requirement that the energy density of GWs should not be too large, as it would alter the Hubble expansion rate and disrupt the successful BBN. The recent joint CMB+BBN analysis implies $\Omega_{\rm GW}h^{2}<1.3\times 10^{-6}$ at $2\sigma$ for $f>2\times 10^{-11}$z Yeh et al. (2022).
@@
This highlights the interesting fact that certain parameter regions can be further constrained by GW constraints. In Fig. 5, the parameter space allowed by the BBN bound $\Omega_{\rm GW}h^{2}<1.3\times 10^{-6}$ is shaded with orange. As indicated in Fig 4, the parameter space predicting too low reheating temperature (larger e-folding number, larger $n_{s}$) is excluded as it predicts too large GW spectrum at high frequencies. Furthermore, the red shaded region indicates where the parameters predict a reasonable reheating temperature. A too low reheating temperature ($T_{\rm reh}\lesssim 10^{-3}$eV) disrupts the successful BBN, while an excessively high reheating temperature ($T_{\rm reh}\sim 4.0\times 10^{15}$eV) implies instant reheating [^4] Note that in the context of supergravity, considerations on the overproduction of gravitinos can further constrain the reheating temperature to $T_{\rm reh}\lesssim 10^{8}-10^{9}$eV Kawasaki et al. (2006); Dimopoulos and Owen (2017) or even down to $10^{6}$eV Kohri et al. (2006).. This temperature corresponds to the e-folding number of $N_{\ast}\sim 56$ and when the e-folding is lower than it we need a phase with $w<1/3$ such as a early matter domination, which requires extra physics to achieve it in the picture of $\alpha$-attractor model. Therefore, the overlap of the red and orange shaded areas represents the allowed region, enabling us to reject more parameter space of the model. We also find that the CMB and LSS constraints indicate a slight preference for lower reheating temperatures; however, higher reheating temperatures are still allowed at the $2\sigma$ level.
@@ VI Conclusions
In this work, we explored the compatibility of a specific class of $\alpha$-attractor models when using the latest observational data, motivated by recent implications of evolving DE equation of state parameter from DESI. The $\alpha$-attractor models present a compelling and unified framework, motivated by high-energy physics and supergravity, as an alternative to traditional single-field slow-roll inflation. These models connect early- and late-time physics, offering a viable and physically motivated alternative to other phenomenological models.
@@
We also considered a connection of the $\alpha$-attractor models with GWs by comparing the spectrum of inflationary GWs predicted in the $\alpha$-attractor model with early kination phase; see Fig. 4, compared with sensitivities of current/future GW experiments. We find the interesting fact that certain parameter regions can be further constrained by the current BBN and CMB constraints on the GW amplitude, see Fig. 5. These constraints could be strengthened by future CMB B-mode measurements as well as constraints by interferometer experiments. Additionally, by considering an appropriate range of reheating temperatures, we demonstrated that the parameter space can be further tightened. However, the analysis we performed here is only indicative, thus we leave a more complete study for future work.
@@ Acknowledgements
The authors thank Nils Schöneberg for useful discussions and they acknowledge the use of the Finis Terrae III supercomputer, which is part of the Centro de Supercomputacion de Galicia (CESGA) and is funded by the Ministry of Science and Innovation, Xunta de Galicia and ERDF (European Regional Development Fund). The authors also acknowledge support from the research project PID2021-123012NB-C43 and the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033. GA’s and SK’s research is supported by the Spanish Attraccion de Talento contract no. 2019-T1/TIC-13177 and 2023-5A/TIC-28945 granted by the Comunidad de Madrid. SK is partly supported by the I+D grant PID2020-118159GA-C42 and PID2023-149018NB-C42 funded by MCIN/AEI/10.13039/501100011033, the Consolidación Investigadora 2022 grant CNS2022-135211, and Japan Society for the Promotion of Science (JSPS) KAKENHI Grant no. 20H01899, 20H05853, and 23H00110. MC acknowledges support from the “Ramón Areces” Foundation through the “Programa de Ayudas Fundación Ramón Areces para la realización de Tesis Doctorales en Ciencias de la Vida y de la Materia 2023”.
@@ Appendix A Quantitative discussion of the $\alpha$-attractor models
The main effect of $\alpha$-attractor models at late times, compared to $\Lambda$CDM, is a change at the background expansion of the Universe causing deviations at a) the CMB spectra at low multipoles via the Integrated Sachs-Wolfe (ISW) effect, and b) at the angular diameter distance, affecting directly the BAO measurements.
@@
Appendix B Comparison with the $w_{0}w_{a}$ model
In this section we also briefly present some constraints on a two-parameter extension of the $\Lambda$CDM model, in terms of the dark energy equation of state parameter, of the form $w(a)=w_{0}+w_{a}\,(1-a)$, where $w_{0}$ and $w_{a}$ are to free parameters. As before, using the CMB+Lensing, SnIa and DESI data we find the constraints as shown in Fig. 10, and compare the results with the ones obtained for the $\alpha$-attractor model.
We observe that the best fit parameters of the $w_{0}w_{a}$ model, are $w_{0}=-0.796_{-0.067}^{+0.061}$ and $w_{a}=-0.972_{-0.270}^{+0.300}$ respectively, which are in very good agreement with the ones reported in Ref. Adame et al. (2024b). The log-evidence for this model is $\ln E_{i}=-2172.89$ and the log Bayes compared to $\Lambda$CDM (see Tab. 3) is $\ln B=13.13$, which shows strong overall preference for the $w_{0}w_{a}$ model.
Figure 10: The $68.3\%$$95.5\%$ confidence contours for the cosmological parameters of the CPL model, which is derived using CMB$+$Lensing, SnIa and DESI data.